# The Wisdom of a Crowd of Brains: A Universal Brain Encoder

### v1

@@ 4 Experiments & Evaluations
In this section we empirically evaluate our Universal-Encoder and its effectiveness for integrating data from multiple subjects/datasets. This section is organized as follows. We first present the datasets and the quantitative metrics used for evaluating the performance of the Universal Encoder. Next (Sec. 4.1), we demonstrate the ability of our Universal-Encoder to jointly train on multiple different subjects who were never exposed to any shared data, thus exploiting the union of all their different training sets. We further show that this exceeds the performance of any individual subject in the cohort. We then show (Sec. 4.2) that old 3T fMRI datasets with limited low-resolution data can be significantly improved by leveraging a new high-quality 7T dataset. Finally (Sec. 4.3), we show that an already trained Universal-Encoder can be easily adapted to new subjects/datasets with minimal amount of new training data, using Transfer-Learning.
Datasets
@@
We experimented with 3 prominent fMRI datasets: (i) The old "vim-1" dataset [23, 39, 40], which features around 1750 train and 120 test grayscale images , and their corresponding 4-Tesla fMRI recordings for 2 subjects. (ii) The “fMRI-on-ImageNet” dataset [41], which comprises 1200 train and 50 test pairs of natural images from ImageNet [42], with 3-Tesla fMRI recordings on 5 subjects. (iii) The "Natural Scenes Dataset" (NSD) [43], a new 7-Tesla dataset with 8 subjects, each having around 9000 unique subject-specific images, and $\sim$1000 images shared across all subjects (which we use as our test set). This results in a total of 73,000 images, all taken from the COCO dataset [42]. A few example images from each of the three datasets are displayed in Figure 1.
Evaluation Metrics
@@
We evaluate our Universal-Encoder (i.e., its ability to correctly predict the fMRI responses of different subjects to a variety of images), using two quantitative measures. Given a set of N Test images with ground-truth fMRI per subject, we first predict the fMRI responses of those N images with our Encoder. We can then compute:
(i) Pearson Correlation (per voxel) – We compute the normalized correlation between the sequence of $N$ predicted fMRI activations of each brain-voxel (on all $N$ Test images), versus the $N$ ground-truth fMRI activations of that voxel on those images.
@@
(ii) Image Retrieval (per image) – For each real fMRI scan in the test set (denoted as “Query”), we aim to retrieve (detect) its corresponding test image which produced it (out of all N test images). We first predict the fMRI from each test-image. We then measure the degree of similarity between these N predicted fMRIs and the real test fMRI scan (via normalized correlation), and sort the N predicted fMRIs accordingly. The test fMRI obtains a Retrieval Ranking score $k$, if the fMRI predicted from its real test image is located in the $k$th place (out of $N$) in the sorted list.
@@
Figure 3 provides a visual (qualitative) example of the Retrieval-Ranking test (for Subject 1 of the NSD dataset). Fig. 3a displays the real and the Encoder-predicted fMRIs for a few sample test images. As can be seen, there is a high similarity between the real fMRI scan and the Encoder-predicted one. Fig. 3b shows the top-5 retrieved images (out of N=1000) for a few example “Query” test-fMRIs. The ground-truth test-image of each Query test-fMRI is displayed in the leftmost column of Fig. 3b. The retrieved images are ranked by the similarity of their Encoder-predicted fMRI to the real “Query” fMRI scan of the test image. As can be seen, there are many distracting (very similar) images among the 1000 test images. Yet, our Universal-Encoder is able to obtain an average retrieval-rank score of 1.85 (out of 1000) for Subject 1 (averaged over all 1000 Query test-fMRIs).
@@
Figure 4: The Wisdom of a Crowd of Brains. By aggregating data from multiple subjects, our Universal-Encoder effectively improves encoding of any individual subject. We compared 3 models: (i) The "Baseline" single-subject encoder of [15], (ii) "Universal Encoder - single subject" – our architecture trained on each subject separately, (iii) "Universal Encoder - multiple subjects" – our model trained on data from 8 subjects. (a) Pearson correlation (per voxel) between predicted & ground-truth fMRI (Median value, 75th & 25th percentiles). (b) Retrieval accuracy: ranking of the GT image within the Top-1 & Top-5 for "Query" fMRI.
@@ 4.1 The Wisdom of  a “Crowd of Brains”
We first demonstrate the Universal-Encoder’s ability to exploit data from multiple subjects, without any shared-data. For this we use the new 7-Tesla NSD dataset. Our Encoder’s train-set comprised the union of all the 8 subject-specific training sets ($\sim$9000 unique images per subject), resulting in a total of $\sim$72,000 pairs of images with their corresponding fMRI scans. Our Encoder’s test-set comprised the $\sim$1,000 shared images (the images that all 8 subjects saw). We compared 3 models in our evaluation: (i) As a baseline, we used the image-to-fMRI encoder of [15], trained separately for each subject on their subject-specific training-set (“Baseline encoder”). (ii) Our Universal-Encoder trained on each subject separately (“Universal Encoder - single subject”), and (iii) Our Universal-Encoder trained on all 8 subjects jointly, using their combined training-sets, and tested on each subject individually (“Universal Encoder - multiple subjects”).
Fig. 4a shows the median Pearson correlation value (along with the $25$th & $75$th percentiles – indicated by a rectangular bar around the median value), computed between the predicted activations and the ground-truth activations for all fMRI voxels. These are computed per subject, for each of the 3 models. Our Universal-Encoder, even when trained on a single subject, performs consistently better than the Baseline-encoder. Moreover, when our Universal-Encoder is trained jointly on the training sets of all subjects, it consistently outperforms all subject-specific models. It obtains notable improvements for both the “best” subjects (e.g., Subject 1 with $\sim$7improvement) and the “worst” subject (Subject 8 with $\sim$15improvement).
Fig. 4b further presents quantitative Retrieval results evaluated per subject for all 3 models. It shows the percent of times the correct image (corresponding to the "Query" fMRI) was ranked $1$st (Top-1) and among the Top-5 retrieved images. The results indicate superior performance of the multi-subject Universal-Encoder compared to the 2 other models in both Top-1 & Top-5. Significance testing of the Universal Encoder’s improvement over other models, as shown in Fig. S8, reveals a maximum p-value of approximately $\sim$3e-12. The above experiments demonstrate that our Universal-Encoder can effectively aggregate data from multiple subjects (who viewed different images), while enhancing the performance of each subject individually.
@@ 4.2 The Wisdom of  a “Crowd of Datasets”
The Universal-Encoder can aggregate fMRI data from multiple datasets, each with own scanning resolution and unique image domain (e.g., B/W vs. color images). This allows training the Universal Encoder jointly on a high-quality (7T) and lower-quality (3T) datasets, thus significantly enhancing the encoding performance of old lower-quality datasets. Fig. 5 demonstrates this by training the encoder on the NSD dataset alongside two low-resolution datasets : VIM1 and “fMRI-on-ImageNet”, and testing the encoding performance on individual subjects within those datasets.
Fig. 5a compares 3 models for the “fMRI-on-ImageNet” dataset: (i) The single-subject encoder of [15] (“Baseline encoder”), (ii) Our Universal-Encoder trained on all subjects within “fMRI-on-ImageNet”(“Universal Encoder - same dataset”), and (iii) Universal-Encoder trained on subjects from both “fMRI-on-ImageNet” and NSD (“Universal Encoder - multi datasets”). Our multi-subject same-dataset encoder (green) outperforms the single-subject baseline model (blue). Adding data from NSD yields even further improvement (red). Median correlation, 75th & 25th percentiles are shown. Fig. 5b shows results for the VIM1 dataset. Here too, adding training data from the high-resolution NSD dataset, significantly enhances encoding performance on VIM1.
Figure 5: The Wisdom of the Crowd of Datasets: Adding data from a high-quality 7-Tesla dataset (NSD) to lower-quality datasets significantly enhances encoding performance in the lower-quality datasets.
@@ 4.3 Transfer-Learning to New Subjects/Datasets
When a new subject/dataset is encountered, it is not necessary to train the universal-Encoder from scratch. Instead, we can add a new subject via quick Transfer-Learning, which is particularly useful when the new subject-specific data is scarce. In our transfer-learning, all weights of the pre-trained Universal-Encoder remain fixed, and only the 256-dimensional Voxel-Embeddings are optimized for the new subject. This allows rapid and effective transfer learning, with little new training data.
@@
To demonstrate the effectiveness of such Transfer-Learning, we pre-train the Universal-Encoder on 6 subjects from the 7-Tesla NSD dataset (Subjects 2-7). We then adapt it via transfer-learning to new subjects (without shared data) – both new subjects within NSD (Subjects 1 & 8), and to subjects in entirely different (older) fMRI datasets (“fMRI-on-ImageNet” & VIM1). Each of the 3 plots in Fig. 6 Compares: (i)  Transfer-Learning of the pre-trained Universal-Encoder to the new subject, with varying numbers of subject-specific training data (purple curve), and (ii) a dedicated subject-specific model, trained from scratch on the subject-specific data only (orange curve). The x-axis represents the number of subject-specific training examples, and the y-axis shows the mean and standard deviation of the median Pearson correlation between the predicted and real fMRI scan from the new subject’s test-set, over 5 runs with different random initialization & data sub-sample.
@@
The transferred Universal-Encoder significantly outperforms any single-subject model on all datasets. Transfer-Learning to Subject 1 within NSD (Fig. 6a) obtains more than 77% improvement for any number of training examples. Moreover, with as little as 100 subject-specific training examples, it already achieves better results than the subject-specific model trained on the entire train-set (of 9000 training examples). A similar gap in performance is achieved for Subject 8. The transferred Universal-Encoder reaches a performance plateau at $\sim$1,000 examples.
@@
Fig. 6b ("fMRI-on-ImageNet") & Fig. 6c (VIM1) demonstrate transfer learning from a new 7-Tesla dataset to older lower-resolution 3-Tesla or 4-Tesla datasets. These were scanned on very different types of images (e.g., VIM1 was scanned on B/W images with a circular black mask), and have much smaller train-sets. Transfer-Learning shows a significant improvement for any number of training data, with an improvement of more than 84% for “fMRI-on-ImageNet” and $\sim$45for VIM1 when using the entire (small) subject-specific train-set. Significance testing shown in appendix Fig. S9.
Figure 6: Transfer-Learning of Universal-Encoder to new subjects/datasets. Pre-trained on NSD, the Universal-Encoder adapts to new subjects with limited data. Transfer-Learning (purple) significantly outperforms single-subject models (orange). The plots show mean and standard deviation of the median Pearson correlation over 5 runs as a function of the number of training examples.
Figure 7: Exploring The Brain: Clustering voxel-embeddings by their proximity in the shared embedding space allows to discover and explore functionality of brain regions. The functional role of each detected cluster of voxels is understood by viewing the images that most strongly activate these clusters (see text for more details).
@@ 5 Exploring the Brain using Voxel-Embeddings
As part of the training, our Universal-Encoder learns the functional role of each brain voxel. It maps functionally-similar brain voxels (whether within the same brain or across different brains) to nearby Voxel-Embeddings in the shared embedding space. This provides a powerful means to explore the human brain and discover new functional regions within it. What facilitates such advanced brain exploration is the enormous number of images that a large “crowd of brains” has collectively been exposed to (which is prohibitive for any single subject). In this section, we demonstrate that these learned voxel-embeddings, which combine data across subjects/datasets, capture semantically meaningful brain functionality and serve as robust voxel representations to explore the brain.
The standard way of dividing the brain into functional regions relies on predefined anatomical regions. This is a very coarse division which overlooks individual differences. We aim to find a division that is functionally-consistent across different subjects, while not being constrained by their anatomical differences. To achieve this, we apply k-means clustering to all brain-voxel-embeddings of multiple “good” subjects (subjects with high prediction accuracy) from the 7-Tesla NSD dataset. Strong clusters often indicate the detection of a meaningful (joint) functionality within the brain.
@@
To uncover the functional roles of such clusters, we examine which image stimuli result in fMRI scans that induce the highest activation levels per cluster (averaged over all voxels within the cluster). Fig. 7a displays an example of the top 4 interesting discovered clusters (displayed on the brain with different colors), along with corresponding top images (images with the highest activation per cluster) for NSD Subjects 1 & 2. Each of these clusters has a clear and distinct functional role, being strongly activated by images of Food, Faces, Text, Outdoor-Scenes, respectively. These results show that our voxel-embeddings capture functional roles rather than individual identities, hence provide a potentially powerful tool to discover shared & unique brain functionalities across different people.
@@
We further explored the ability to detect finer functional granularity within known brain regions. Fig. 7b shows one such example – the detection of functionally-meaningful clusters (sub-regions) within the PPA brain region (an area corresponding to places/scenes). Two clear and distinct sub-regions have emerged from our voxel-embedding clustering – Indoor-Scenes vs. Outdoor-Scenes. This demonstrates the power of our embeddings to uncover new functional regions beyond predefined anatomical boundaries. Additional examples & experiments are found in the appendix (Figs.S10S11).
6 Conclusion & Discussion
@@
This paper presents the first-ever Universal Image-to-fMRI Brain-Encoder, which can integrate data from many different subjects and different fMRI datasets collected over the years. This is facilitated by a new voxel-centric architecture, which learns individual “voxel-embedding” per brain-voxel, via cross-attention with hierarchical image features. In this voxel-centric architecture, the functional role of individual brain voxels naturally emerges, leading to better encoding performance and providing a new tool to explore brain functionality (e.g., what is encoded where in the brain). Moreover, this architecture is indifferent to the number of brain-voxels in each fMRI scan, hence can be applied to any fMRI data.
Our approach could potentially be extended to developing Brain-Encoders for other data modalities (e.g., video, audio, speech). Moreover, the Voxel-Embeddings may potentially provide a means for exploring the existence or lack of brain functionalities in ‘irregular’ brains (such as autistic brains, or brains of visually impaired people). Since a trained Universal-Encoder can adapt to new subjects/datasets with minimal new data via Transfer-Learning, it could practically explore such new domains with only a small number of new subject-specific scans required.

### v2

@@ 4 Experiments & Evaluations
We first present the datasets and the quantitative metrics used for evaluating the performance of the Universal Encoder. Then (Sec. 4.1), we demonstrate the ability of our Universal-Encoder to jointly train on multiple different subjects who were never exposed to any shared data, thus exploiting the union of all their different training sets. We further show that this exceeds the performance of any individual subject in the cohort. We then show (Sec. 4.2) that old fMRI datasets with limited low-resolution can be significantly improved by leveraging a new high-quality 7T dataset. Finally (Sec. 4.3), we show that an already trained Universal-Encoder can be easily adapted to new subjects/datasets with minimal amount of new training data, using Transfer-Learning.
Data & Datasets:
We use publicly available fMRI datasets, which contain pairs of images and their corresponding fMRI scans. These capture the brain activity of different subjects viewing images. Each fMRI scan provides measurements of neural activity in small volumes of the cortex (the “brain-voxels” in this paper). After pre-processing (performed by the groups who collected these data), each brain voxel is represented by a single scalar value (the average neural activation within that brain voxel).
@@
We experimented with 3 prominent fMRI datasets: (i) The old “vim1” dataset (Kay et al., 2008; Naselaris et al., 2009; Kay et al., 2011), which has 1750 train and 120 test grayscale images, and their corresponding 4-Tesla fMRI scans for 2 subjects. (ii) The “fMRI-on-ImageNet” scans (Horikawa & Kamitani, 2017b) comprising 1200 train and 50 test pairs of natural images from ImageNet, with 3-Tesla fMRI recordings on 5 subjects. (iii) The ”Natural Scenes Dataset” (NSD) (Allen et al., 2022), a new 7-Tesla dataset with 8 subjects, each having around 9000 unique subject-specific images, and $\sim$1000 images shared across all subjects (which we use as our test set). This results in a total of 73,000 images, all taken from the COCO dataset (Deng et al., 2009). A few example images from each of the 3 datasets are displayed in Fig. 1. A subset of visually sensitive brain voxels was either provided by the dataset or selected based on SNR (see Appendix A.1)
Evaluation Metrics:
@@
We evaluate our Universal-Encoder (its ability to correctly predict the fMRI responses of different subjects to a variety of images), using 2 quantitative measures. Given a set of $N$ Test images with ground-truth fMRI per subject, we first predict the fMRI responses of those N images with our Encoder. We then compute:
@@
(i) Pearson Correlation (per voxel) – Normalized correlation between the sequence of $N$ predicted fMRI activations of each brain-voxel (on all $N$ Test images), versus the $N$ ground-truth fMRI activations of that voxel on those images. (ii) Image Retrieval (per image) – For each real fMRI scan in the test set (denoted “Query”), we aim to retrieve (detect) its corresponding Test-image which produced it out of a set of N images (the Test-image and N-1 random distractors). To do so, we first predict the fMRIs of all N images in the set (using our Universal-Encoder). We then search for the Nearest-Neighbor (NN) of the real test-fMRI (Query) among the set of N predicted fMRIs (using Pearson correlation). If the fMRI predicted from the Test-image is retrieved as the $1^{st}$ NN, it obtains a “Rank-1” score. If it is retrieved as the $k^{th}$ NN, it obtains a “Rank-k” score. The reported Top-1 accuracy is the percent of test-fMRIs which obtained a Rank-1 score. Top-5 accuracy is the percent of test-fMRIs which obtained a score $\leq$ Rank-5 (i.e., the corresponding test image was among the first 5 NNs).
@@
Fig. 3 shows a qualitative visual example of the Retrieval-Ranking test (for Subject1 of the NSD dataset). Fig. 3a displays the real and the Encoder-predicted fMRIs for a few sample test images. As can be seen, there is high similarity between the real fMRI scan and the Encoder-predicted one. Fig. 3b shows the top-5 retrieved images (out of N=1000) for a few example “Query” test-fMRIs. The ground-truth test-image of each Query test-fMRI is displayed in the leftmost column of Fig. 3b. The retrieved images are ranked by the similarity of their Encoder-predicted fMRI to the real “Query” fMRI scan of the test image. As can be seen, there are many distracting (very similar) images among the 1000 test images. Yet, our Universal-Encoder is able to obtain an average retrieval-rank score of 1.85 (out of 1000) for Subject 1 (averaged over all 1000 Query test-fMRIs). Appendix Fig. S13 further demonstrates that our image-to-fMRI encoder preserves relevant image information, as evidenced by successful fMRI-to-image reconstructions.
@@
Figure 4: The Wisdom of a Crowd of Brains. By aggregating data from multiple subjects, our Universal-Encoder improves encoding of any individual subject. We compared 3 models: (i) ”Baseline” single-subject encoder of (Gaziv et al., 2022), (ii) ”Universal Encoder - single subject” – our architecture trained on each subject separately, (iii) ”Universal Encoder - multiple subjects” – our model trained on data from 8 subjects. (a) Pearson Correlation (per voxel) between predicted & ground-truth fMRI (Median value, 75th & 25th percentiles). (b) Retrieval Accuracy (Top-1 & Top-5) of the GT image per “Query” fMRI.
@@ 4.1 The Wisdom of  a “Crowd of Brains”
We first demonstrate the Universal-Encoder’s ability to exploit data from multiple subjects, without any shared-data. For this we use the 7-Tesla NSD dataset. We train on union of all 8 subject-specific images ($\sim$72,000 non-shared data), and test on the $\sim$1,000 shared images (the images that all 8 subjects saw). We compared 3 models in our evaluation: (i) As a baseline, we used the image-to-fMRI encoder of (Gaziv et al., 2022), trained separately for each subject on their subject-specific training-set (“Baseline encoder”). (ii) Our Universal-Encoder trained on each subject separately (“Universal Encoder - single subject”), and (iii) Our Universal-Encoder trained on all 8 subjects jointly, using their combined training-sets, and tested on each subject individually (“Universal Encoder - multiple subjects”). A comparison to two other prominent encoder models (Wang et al., 2023; Takagi & Nishimoto, 2023) (which have lower performance), is further provided in the appendix Fig. S8.
Fig. 4a shows the median Pearson correlation value (along with the $25$th & $75$th percentiles), computed between the predicted activations and the ground-truth activations for all fMRI voxels. These are computed per subject, for each of the 3 models. Our Universal-Encoder, even when trained on a single subject, performs consistently better than the Baseline-encoder. Moreover, when our Universal-Encoder is trained jointly on the training sets of all subjects, it consistently outperforms all subject-specific models. It obtains notable improvements for both the “best” subjects (Subject1: $\sim$7improvement) and the “worst” subject (Subject8: $\sim$15improvement). Statistical significance of the performance gap is assessed in Appendix C, confirming significant superiority of the Universal-Encoder. Per-region results are further presented in Fig. S10.
Figure 5: The Wisdom of the Crowd of Datasets: Using data from a high-quality 7T dataset (NSD) significantly enhances the encoding performance in lower-quality (3T & 4T) datasets.
Fig. 4b presents quantitative Retrieval results evaluated per subject for all 3 models. It shows the percent of times the correct image (corresponding to the ”Query” fMRI) was ranked $1$st (Top-1) and among the Top-5 retrieved images. The results indicate superior performance of the multi-subject Universal-Encoder compared to the 2 other models in both Top-1 & Top-5. Statistical significance of the Universal Encoder’s improvement over other models is verified in Appendix C & Table.S2. Our experiments demonstrate the model’s ability to effectively aggregate data from multiple subjects (who viewed different images), while enhancing the performance of each subject individually.
@@ 4.2 The Wisdom of  a “Crowd of Datasets”
The Universal-Encoder can combine fMRI data from multiple datasets, each with its own scanning resolution and unique image domain (e.g., B/W vs. color images). This allows training the Universal Encoder jointly on high-quality (7T) and lower-quality (3T) datasets, thus significantly enhancing the encoding performance of old lower-quality datasets. Fig. 5 demonstrates this by training the encoder on the NSD dataset alongside two low-resolution datasets: VIM1 and “fMRI-on-ImageNet”, and testing the encoding performance on individual subjects within those datasets.
Fig. 5a compares 3 models for the “fMRI-on-ImageNet” dataset: (i) Single-subject encoder of (Gaziv et al., 2022) (“Baseline encoder”), (ii) Our Universal-Encoder trained on all subjects in “fMRI-on-ImageNet” (“Universal Encoder - same dataset”), and (iii) Universal-Encoder trained on subjects from both “fMRI-on-ImageNet” and NSD (“Universal Encoder - multi datasets”). Our multi-subject same-dataset encoder (green) outperforms the single-subject baseline model (blue). Adding data from NSD yields even further improvement (red). Median correlation, 75th & 25th percentiles are shown. Fig. 5b shows results for VIM1 dataset. Here too, adding training data from NSD significantly enhances encoding performance on VIM1.
Figure 6: Transfer-Learning to new subjects/datasets. Pre-trained on NSD, the Universal-Encoder adapts to new subjects with few data. Transfer-Learning (purple) significantly outperforms single-subject models (orange). See text for details.
@@ 4.3 Transfer-Learning to New Subjects/Datasets
When a new subject/dataset is encountered, it is not necessary to train the universal-Encoder from scratch. Instead, we can add a new subject via quick Transfer-Learning, which is particularly useful when the new subject-specific data is scarce. In our transfer-learning, all weights of the pre-trained Universal-Encoder remain fixed, and only the 256-dimensional Voxel-Embeddings of the new subject are optimized. This allows rapid and effective transfer learning, with very little new training data.
@@
To demonstrate this, we pre-train the Universal-Encoder on 6 subjects from the 7T NSD dataset (Subjects 2-7), freeze it, and then apply Transfer-Learning to new subjects (with NO shared data): Subject1 in NSD, and subjects in entirely different (older) fMRI datasets (“fMRI-on-ImageNet” (subject3) & VIM1 (subject1)). Each plot in Fig. 6 compares: (i) Transfer-Learning of the pre-trained Universal-Encoder to the new subject, with varying numbers of subject-specific training data (purple curve), and (ii) a dedicated subject-specific model, trained from scratch on the subject-specific data only (orange curve). The x-axis shows the number of subject-specific training examples, and the y-axis shows mean and standard deviation of the median Pearson correlation between the predicted and real fMRI scan from the new subject’s test-set (over 5 runs with different random initialization & data sub-sampling).
@@
The transferred Universal-Encoder significantly outperforms any single-subject model on all datasets. Transfer-Learning to Subject1 within NSD (Fig. 6a) obtains $\geq$77improvement for any number of training examples. Moreover, as few as 100 subject-specific training examples suffice to obtain better results than a subject-specific model trained on the entire train-set (9000 examples). A similar gap in performance is achieved for Subject8.
Figure 7: Brain Exploration: Clustering voxel-embeddings in the Embedding-Space leads to discovery of functional brain regions. The functional role of each embedding cluster is unveiled through the images that give highest activation to each cluster (see text).
@@
Fig. 6b (”fMRI-on-ImageNet”) & Fig. 6c (VIM1) demonstrate transfer learning from a new 7-Tesla dataset to older lower-resolution 3T or 4T datasets. These were scanned on very different types of images (e.g., VIM1 was scanned on B/W images with a circular black mask), and have much smaller train-sets. Transfer-Learning shows a significant improvement for any number of training data, with an improvement $\geq$84for ‘fMRI-on-ImageNet’ & $\sim$45for VIM1, using the entire (small) subject-specific train-set (significance testing over model initialization in Table. S3).
@@ 5 Exploring the Brain using Voxel-Embeddings
As part of training, our Universal-Encoder aims to learn the functional-role of each brain voxel. It maps functionally-similar brain voxels (within the same brain & across different brains) to similar voxel-embeddings. This provides a potentially powerful means to explore the human brain and discover new functional regions. What facilitates such advanced brain exploration is the enormous number of images that a large “crowd of brains” has collectively been exposed to (which is prohibitive for any single subject). In this section, we demonstrate initial promising findings, suggesting that our learned voxel-embeddings capture semantically-meaningful brain functionalities, and may potentially serve as a new powerful means to explore the human brain.
Brain-parcellations which define functional brain regions do exist. Although functionally-driven, these maps are typically integrated across many subjects via pure anatomical alignment. This results in a single coarse functional brain division, which is shared by all subjects, and overlooks individual differences. We aim to find a division that is functionally-consistent across different subjects, without requiring shared anatomical mapping, thus allowing for individual anatomical (and functional) differences.
To automatically detect functional regions shared across different brains, we performed the following experiment: We applied k-means clustering to all voxel-embeddings of multiple “good” subjects (subjects with high prediction accuracy) from the 7T NSD dataset. Large clusters of voxel-embeddings indicate the detection of some joined functionality (learned independently by many different brain-voxels). Since high similarity of independently-learned voxel-embeddings is unlikely to occur at random, this indicates that they most likely encode something meaningful.
@@
To unveil the functional roles of each detected embedding cluster, we examine which images induce the highest fMRI activation per cluster (averaged over all voxels within the cluster). Fig. 7a displays an example of the top 4 (out of 20) interesting discovered clusters (displayed on the brain with different colors), along with corresponding top images (with highest activation per cluster) for NSD Subjects 1&2. Each cluster in Fig. 7a shows a distinct functional-role, being strongly activated by images of Food, Faces, Text, Outdoor-Scenes, respectively. It is interesting to note that: (i) These shared functional regions across brains automatically surfaced out, although the 2 subjects viewed different images; (ii) While the shared regions across the 2 subjects are in nearby anatomical brain locations, they are not anatomically aligned! These results show that our voxel-embeddings capture functional roles rather than individual identities, providing a potentially powerful tool to discover shared&unique brain functionalities across different people. Additional detected voxel-embedding clusters can be viewed in Appendix Figs. S14,S15.
@@
We further explored the ability to detect finer-grained functionality within known brain regions. Fig. 7b shows such an example - the detection of functionally-meaningful clusters (sub-regions) within the PPA brain region (an area corresponding to places/scenes). Two clear and distinct sub-regions have emerged on their own from our voxel-embedding clustering – Indoor-Scenes vs. Outdoor-Scenes. This demonstrates the potential power of our voxel-embeddings to uncover new functional regions beyond predefined anatomical boundaries. Additional visualizations are found in the appendix: e.g., fine-grained clustering of EBA brain region (Fig. S16), T-SNE visualization for voxel-embeddings of all subjects (Figs. S17 & S18).
Conclusion & Discussion:
@@
This paper presents the first-ever Universal Image-to-fMRI Brain-Encoder, which can integrate data from many different subjects and different fMRI datasets collected over the years. This is facilitated by a new voxel-centric architecture, which learns individual “voxel-embedding” per brain-voxel, via cross-attention with hierarchical image features. In this voxel-centric model, the functional role of individual brain voxels naturally emerges, leading to better encoding performance and providing a new potential tool to explore brain functionality (i.e., what is encoded where in the brain).
Acknowledgments
This research was funded by the European Union (ERC grant No. 101142115).